Find the slope of any line that is perpendicular to the line whose equation is 2 x - 3 y + 4 = 0
Answer by Alan3354(69443)

Find the slope of any line that is perpendicular to the line whose equation is 2 x - 3 y + 4 = 0
-------------
3x + 2y = k (k = any number)
---------
Swap the coefficients and change the sign.